OpenStudy (anonymous):

Solve for variable r: 1/r +1/t = 1/v

OpenStudy (solomonzelman):

\[\frac{1}{r}+\frac{1}{t}=\frac{1}{v}\]\[\frac{1}{r}=\frac{1}{v}-\frac{1}{t}\]\[\frac{1}{r}=\frac{t-v}{tv}\]\[r(t-v)=tv\]\[r=\frac{tv}{(t-v)}\]

OpenStudy (anonymous):

How do you get to the third step exactly?